OpenStudy (anonymous):

completing the square x^2+x+1=0

OpenStudy (anonymous):

i think ur question is x^2+2x+1=0 (x+1)(x+1) i think its like this !

OpenStudy (anonymous):

not that

myininaya (myininaya):

\[x^2+x=-1\] \[x^2+1\cdot x+(\frac{1}{2})^2=-1+(\frac{1}{2})^2\] \[(x+\frac{1}{2})^2=-1+\frac{1}{4}\] \[(x+\frac{1}{2})^2=\frac{-3}{4}\] \[x+\frac{1}{2}=\pm \sqrt{\frac{-3}{4}}\] \[x=-\frac{1}{2} \pm \sqrt{\frac{-3}{4}}\] \[x=-\frac{-1}{2} \pm i \frac{\sqrt{3}}{2}\]

OpenStudy (lgbasallote):

the steps: x^2 + x + 1/4 = -1 + 1/4 (x+1/2)^2 = -3/4 x + 1/2 = (sqrt of -3 )over 2 x = [(i sqrt of 3) - 1]/2